### v1

# Evidence for Primordial Alignment: Insights from Stellar Obliquity Measurements for Compact Sub-Saturn Systems

Abstract
Despite decades of effort, the mechanisms by which the spin axis of a star and the orbital axes of its planets become misaligned remain elusive. Particularly, it is of great interest whether the large spin-orbit misalignments observed are driven primarily by high-eccentricity migration — expected to have occurred for short-period, isolated planets — or reflect a more universal process that operates across systems with a variety of present-day architectures. Compact multi-planet systems offer a unique opportunity to differentiate between these competing hypotheses, as their tightly-packed configurations preclude violent dynamical histories, including high-eccentricity migration, allowing them to trace the primordial disk plane. In this context, we report measurements of the sky-projected stellar obliquity () via the Rossiter-McLaughlin effect for two sub-Saturns in multiple-transiting systems: TOI-5126 b () and TOI-5398 b (). Both are spin-orbit aligned, joining a fast-growing group of just three other compact sub-Saturn systems, all of which exhibit spin-orbit alignment. Our results strongly suggest that sub-Saturn systems are primordially aligned and become misaligned largely in the post-disk phase through violent dynamical interactions inherent to eccentric migration, as appears to be the case increasingly for other exoplanet populations.
1 Introduction
Stellar obliquity—i.e., the angle between the spin axis of the host star and the net orbit normal axis of the planets—is a powerful probe of a planetary system’s formation history. While the precise mechanisms driving spin-orbit misalignment are not yet clear, it is important to distinguish whether misalignment is largely a product of high-eccentricity migration, which should be confined to short-period isolated planets111Planets that are single or have only distant planetary companion(s) (e.g., hot Jupiters; see Dawson & Johnson 2018 and references therein), or some universal process that operates across planetary systems with a variety of architectures, e.g., magnetic warping (Lai et al., 2011; Romanova et al., 2013), interactions with stellar companions (Lubow & Ogilvie, 2000; Batygin, 2012), stellar flybys (Hao et al., 2013), or stellar gravity waves (Rogers et al., 2012, 2013).
Compact multi-planet systems may be critical to differentiate between these two scenarios, as their tight orbital configurations are inconsistent with high-eccentricity migration induced by planet-planet scattering (Rasio & Ford, 1996; Chatterjee et al., 2008), Lidov-Kozai cycling (Wu & Murray, 2003; Fabrycky & Tremaine, 2007; Naoz, 2016), or secular interactions (Naoz et al., 2011; Wu & Lithwick, 2011; Petrovich, 2015) — all of which tend to disrupt the orbits of nearby planetary neighbors (Mustill et al., 2015). As such, if high-eccentricity migration is the primary driver of misalignment, compact multi-planet systems will retain their primordially aligned configurations. On the other hand, if universal misalignment processes dominate, compact systems may be misaligned at a rate comparable to isolated systems.
The great majority of spin-orbit angle measurements have been made for close-in Jupiters, particularly hot Jupiters (see recent review by Albrecht et al. 2022), which rarely host nearby companions (Steffen et al., 2012; Huang et al., 2016; Hord et al., 2021; Wu et al., 2023). Therefore, it is useful to expand the census of stellar obliquities to other types of exoplanets. Sub-Saturns, broadly defined as giant planets with masses ranging between that of Neptune (17 M) and Saturn (95 M), represent a particularly interesting population for such studies. The core accretion model implies that they are ‘failed gas giants’ (Pollack et al., 1996), suggesting they have analogous origins to Jupiters (Dong et al., 2018; Lee, 2019; Hallatt & Lee, 2022). However, sub-Saturns are less amenable to detection than Jupiters and are relatively understudied as an exoplanet population, so few constraints on their formation histories exist. As of this writing, only about two dozen sub-Saturns have had their spin-orbit angle measured, demonstrating that these planets can take on a range of values from near-perfect alignment to almost completely retrograde. Yet, only three of these measurements have been made for sub-Saturns in compact multi-planet systems: Kepler-9 b (Wang et al., 2018), AU Mic b (Hirano et al., 2020a), and WASP-148 b (Wang et al., 2022). While this sample is limited in number, all three systems show evidence of spin-orbit alignment and thus provide tentative support for primordial alignment.
The Transiting Exoplanets Survey Satellite (TESS) mission (Ricker et al., 2015), now well into its Extended Mission phase, has produced a growing sample of sub-Saturns around bright stars. Of particular interest to this work, TESS is responsible for the discovery of at least two sub-Saturns in compact multi-planet systems: TOI-5126 b (Fairnington et al., 2023), and TOI-5398 b (Mantovan et al., 2022, 2024). Given the brightness of their host stars and advances in Extreme Precision Radial Velocity (EPRV), these systems are amenable to stellar obliquity measurements.
[…]
This paper is organized as follows. In Section 2, we outline our observational methods. In Section 3, we describe our determination of stellar parameters. In Section 4, we detail our modeling of the RM signals and, subsequently, the stellar obliquities. In Section 5, we quantify the significance of our results and discuss their implications for the origins of spin-orbit misalignment and the formation histories of sub-Saturns.
2 Observations
The RM effect of TOI-5126 and TOI-5398 were observed using the WIYN/NEID spectrograph in the High Resolution (HR) mode (, Halverson et al. 2016; Schwab et al. 2016) on February 28, 2023 and April 16, 2023, respectively. NEID is a fiber-fed (Kanodia et al., 2018), actively environmentally stabilized spectrograph (Stefansson et al., 2016; Robertson et al., 2019) with a wavelength coverage of 380 nm to 930 nm. For TOI-5126 b, we obtain 20 radial velocity (RV) measurements in HR mode with 1000-second exposures from 02:20-07:48 UT, covering its full transit. These observations occurred under atmospheric conditions with a seeing range of 1.2-2.8 (median 1.75) and an airmass range of . At a wavelength of , the NEID spectrograph achieved a signal-to-noise ratio (SNR) of 56 pixel. For TOI-5398 b, nearly the full transit (excluding ingress) was covered by 14 RV measurements obtained in HR mode with 1250-second exposures from 02:55-09:04 UT. The conditions featured a seeing range of 0.7-1.5 (median 1.0) and an airmass range of . At the same wavelength, the spectrograph’s SNR was 60 pixel. An additional 14 out-of-transit NEID observations of TOI-5398 b were obtained between April 8, 2022, and December 27, 2023, with 300-second exposure times.
The NEID spectra were analyzed using version 1.3.0 of the NEID Data Reduction Pipeline (NEID-DRP)222Detailed information is available at: https://neid.ipac.caltech.edu/docs/NEID-DRP/, and radial velocities were derived via the cross-correlation function (CCF) method within NEID-DRP. We extracted the resulting barycentric-corrected radial velocities for re-weighted orders (CCFRVMOD) from the NExScI NEID Archive333https://neid.ipac.caltech.edu/. The RV data gathered from NEID for this study are illustrated in Figure 1 (in baseline-subtracted form) and available through the Data Behind the Figure program (see figure caption).
3 Stellar parameters
3.1 Synthetic Spectral Fit By iSpec
We adopt the co-added NEID spectra of TOI-5126 and TOI-5398 to determine their respective stellar atmospheric parameters, including stellar effective temperature ( ), surface gravity ( ), metallicity ( ), and projected rotational velocity ( ). We use the synthetic spectral fitting technique provided by the Python package iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019) to measure these parameters.
We employ the SPECTRUM radiative transfer code (Gray & Corbally, 1994), the MARCS atmosphere model (Gustafsson et al., 2008), and the sixth version of the GES atomic line list (Heiter et al., 2021), all integrated within iSpec, to create a synthetic model for the combined NEID spectra (SNR of 206 for TOI-5126; SNR of 176 for TOI-5398). In our fitting process, we treat micro-turbulent velocities as a variable, allowing us to accurately represent the small-scale turbulent motions in the stellar atmosphere. Conversely, we determine macro-turbulent velocities using an empirical relationship (Doyle et al., 2014) that leverages established correlations with various stellar attributes. We select specific spectral regions to streamline the fitting process, focusing on the wings of the H, H, and Mg I triplet lines, which are indicative of and , as well as Fe I and Fe II lines, which are crucial for constraining and . We then utilize the Levenberg-Marquardt nonlinear least-squares fitting algorithm (Markwardt, 2009) to refine our spectroscopic parameters by iteratively minimizing the value between the synthetic and observed spectra. The final spectroscopic parameters are detailed in Table 1.
3.2 SED+MIST Fit By EXOFASTv2
To ascertain additional stellar parameters, such as stellar mass () and radius (), we utilize the MESA Isochrones & Stellar Tracks (MIST) model (Choi et al., 2016; Dotter, 2016) in combination with a spectral energy distribution (SED) fit. Photometry for the SED fit was compiled from various catalogs, including 2MASS (Cutri et al., 2003), WISE (Cutri et al., 2021), TESS (Ricker et al., 2015), and Gaia DR3 (Gaia Collaboration et al., 2023). We apply Gaussian priors based on our synthetic spectral fit to and , along with the parallax from Gaia DR3, and an upper limit for the -band extinction from Schlafly & Finkbeiner (2011). To accommodate the approximate 2.4% systematic uncertainty floor in , as suggested by Tayar et al. (2022), we increase the uncertainties of to 150 K for both TOI-5126 and TOI-5398.
As mentioned in Fairnington et al. (2023) and Mantovan et al. (2024) respectively, both TOI-5126 and TOI-5398 exhibit strong stellar rotation signals, providing additional stringent constraints on their stellar ages. We adopt Equation 7 from Barnes (2007) to calculate the gyrochronological ages for these stars based on the stellar rotation periods derived from our work (see Section 4). The resulting gyrochronological ages for TOI-5126 and TOI-5398 are Gyr and Gyr, respectively. Subsequently, we adopt these ages and their 3 uncertainties as priors for the stellar ages in our SED fit.
We perform the SED fitting using the Differential Evolution Markov Chain Monte Carlo (DEMCMC) technique, integrated within EXOFASTv2 (Eastman, 2017; Eastman et al., 2019), from which we obtain uncertainty estimates. The DEMCMC procedure was considered converged when the Gelman-Rubin diagnostic (; Gelman & Rubin, 1992) fell below 1.01 and the count of independent draws surpassed 1000. Our final stellar parameters are listed in Table 1.
| Description | Priors | TOI-5126 | Priors | TOI-5398 | |
| Stellar Parameters: | |||||
| Synthetic spectral fit: | |||||
| Effective temperature (K) | - | - | |||
| Metallicity (dex) | - | - | |||
| Surface gravity (log(cm/s)) | - | - | |||
| SED+MIST fit (adopted): | |||||
| Stellar mass () | - | - | |||
| Stellar radius () | - | - | |||
| Effective temperature (K) | |||||
| Metallicity (dex) | |||||
| Surface gravity (log(cm/s)) | - | - | |||
| Age | Age (Gyr) | ||||
| Parallax (mas) | |||||
| d | Distance (pc) | - | - | ||
| (iSpec) | Projected stellar | - | - | ||
| rotational velocity (km/s) | |||||
| Rossiter-McLaughlin Parameters: | |||||
| Sky-projected spin-orbit angle (deg) | |||||
| Projected stellar | |||||
| rotational velocity (km/s) | |||||
| Micro-turbulent velocity (km/s) | |||||
| Macro-turbulent velocity (km/s) | |||||
| Planetary Parameters: | |||||
| Planet-to-star radius ratio | |||||
| Ratio of the sum of star and | |||||
| planet radii to semi-major axis | |||||
| Cosine of inclination | |||||
| Mid-transit time - 2459600 () | |||||
| Orbital period (days) | |||||
| Radial velocity semi-amplitude (m/s) | |||||
| Linear limb-darkening coefficient for NEID | |||||
| Quadratic limb-darkening coefficient for NEID | |||||
| Derived Parameters: | |||||
| Planetary mass () | - | - | |||
| Planetary radius () | - | - | |||
| Semi-major axis scaled by stellar radius | - | - | |||
| Semi-major axis (AU) | - | - | |||
| Inclination (deg) | - | - | |||
| Total transit duration (hours) | - | - | |||
| Linear limb-darkening coefficient for NEID | - | - | |||
| Quadratic limb-darkening coefficient for NEID | - | - | |||
| GP Parameters: | |||||
| Amplitude scaling factor | - | - | |||
| Balance factor | - | - | |||
| Coherence timescale | - | - | |||
| Stellar rotation period | - | - | |||
| Amplitude scaling factor | - | - | |||
| Balance factor | - | - | |||
| Coherence timescale | - | - | |||
| Stellar rotation period | - | - | |||
4 Obliquity Modeling
We determine the sky-projected spin-orbit angle for TOI-5126 b and TOI-5398 b by jointly fitting their two-minute cadence TESS lightcurves and radial velocities from NEID using a modified version of allesfitter package (Günther & Daylan, 2021), along with ancillary ground-based data available from their discovery papers. The modified version of allesfitter incorporates the Hirano et al. (2011) RM model implemented in tracit (Hjorth et al., 2021; Knudstrup & Albrecht, 2022). We exclude the transit data for the non-sub-Saturn planet in both systems but model their RVs (TOI-5126 c at and TOI-5398 c at ).
For TOI-5126 b, we modeled eight full transits and one partial transit from TESS Sector 45, four full transits and one partial transit from Sector 46, and four full transits from Sector 48. Within allesfitter, we simultaneously fit the TESS and NEID data together with all photometric data utilized in TOI-5126 b’s discovery paper (Fairnington et al., 2023), including observations from the CHaracterising ExOplanets Satellite (CHEOPS, Benz et al. 2021) as well as ground-based telescopes at McDonald Observatory in Fort Davis, TX and at the Cerro Tololo Inter-American Observatory (CTIO) in Cerro Tololo, Chile, which are both part of the Las Cumbres Observatory Global Network (LCOGT; Brown et al. 2013). We do not make use of the out-of-transit radial velocities presented in Fairnington et al. (2023) due to their relatively large scatter.
For TOI-5398 b, we model two full transits from TESS Sector 48, as well as photometry from the LCOGT telescopes at McDonald Observatory and Teide Observatory and from the KeplerCam at the Fred Lawrence Whipple Observatory (FLWO) on Mount Hopkins, Arizona (all sourced from Mantovan et al. 2024). We additionally incorporate the out-of-transit radial velocities originally presented in Mantovan et al. (2024), which were obtained from the High Accuracy Radial velocity Planet Searcher (HARPS-N, Cosentino et al. 2012) at the Italian Telescopio Nazionale Galileo. To account for strong stellar activity in the RVs, we apply a rotational Gaussian Process (GP) regression kernel formulated by Foreman-Mackey et al. (2017),
where represents the period of stellar rotation, denotes the coherence timescale, signifies the interval between two successive data points, is a scaling hyperparameter that adjusts the amplitude, and serves as the balance parameter for the periodic and nonperiodic components of the GP kernel.
For both TOI-5126 b and TOI-5398 b, we adopt priors but utilize the results reported in the literature (Fairnington et al. 2023 and Mantovan et al. 2024, respectively) as our initial guesses for , , , , , and (see Table 1 for parameter definitions). We fixed eccentricity () and the argument of periastron () to 0 due to the near-zero eccentricity reported by the discovery papers of TOI-5126 b and TOI-5398 b. The two transformed quadratic limb-darkening coefficients and are initialized at 0.5 for each instrument. As an initial guess for , for which we apply Gaussian priors, we adopt the values derived from our synthetic spectral fit. Finally, we initialize at for both systems. All parameters (except and ) are allowed to vary during the fitting process. Note that, to account for the short-term overnight instrumental systematics and stellar variability in the RM fit, we employ a quadratic function.
For the fits of both systems, we sample the posterior distributions for all modeled parameters using an affine-invariant Markov Chain Monte Carlo (MCMC) grounded in emcee (Foreman-Mackey et al., 2013a) with 100 walkers. Best-fit parameters were obtained after 200,000 accepted steps per walker were reached (note that each ran 10,000 burner steps); our final MCMC results and associated uncertainties are summarized in Table 1. Additionally, all Markov Chains reached their autocorrelation lengths, indicating convergence.
We find good agreement ( difference) with all stellar, planetary, and derived parameters reported in the discovery papers of both systems (Fairnington et al. 2023 for TOI-5126 b and Mantovan et al. 2024 for TOI-5398 b), except (2.5 discrepant) for TOI-5126, which may be due to tentative transit timing variations (see Fairnington et al. 2023). Critically, our RM fits show that both sub-Saturns are consistent with spin-orbit alignment, with TOI-5126 b having a best-fit and TOI-5398 b having a best-fit . Figure 1 reveals the best-fit joint models and associated residuals for both systems, while Table 1 contains all fitted and derived parameters.
It is useful to additionally measure the 3D stellar obliquity , which yields a system’s true spin-orbit configuration. To estimate , we first derive the stellar inclination and stellar equatorial velocity for both TOI-5126 and TOI-5398 by applying the Bayesian inference method presented by Masuda & Winn (2020) and Hjorth et al. (2021) to our values for stellar radius , stellar rotation period , and . We compute the stellar rotation periods via the autocorrelation function (ACF) implemented in SpinSpotter (Holcomb et al., 2022), finding days and days for TOI-5126 and TOI-5398, respectively. The resulting stellar rotation periods are consistent with those reported by Fairnington et al. (2023) for TOI-5126 ( days) and Mantovan et al. (2024) ( days) for TOI-5398. From our Bayesian analysis, we find and for TOI-5126 and and for TOI-5398, and compute the true 3D stellar obliquities as (see Eqn. 9 of Fabrycky & Winn 2009):
where is the inclination angle of the planet. For TOI-5126 b, we find a true obliquity of . Similarly, we derive for TOI-5398 b, demonstrating that both systems are consistent with alignment or near-alignment. Note that, in our derivation of , we adopt an uncertainty floor of 10% in the stellar rotation periods due to the latitudinal differential rotation as suggested in Epstein & Pinsonneault (2014) and Aigrain et al. (2015) and a systematic uncertainty floor of 4.2% in stellar radii reported by Tayar et al. (2022)
5 Discussion
While sub-Saturns broadly appear to boast a wide range of stellar obliquities, the few previously confirmed to be in compact multi-planet systems are exclusively aligned. Our RM measurements for two of such recently-confirmed sub-Saturns, TOI-5126 b and TOI-5398 b, continue this trend, as both are consistent with alignment. We show the updated stellar obliquity distribution for the sub-Saturn population as a function of stellar effective temperature in Figure 2. To construct this sub-Saturn sample, we combine the catalogs of Albrecht et al. (2022) and TEPCat444https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html (Southworth, 2011) accessed on March 7, 2024. We consider only RM effect measurements, which exhibit near-uniform sensitivity to the full range of possible values (compared to nonuniform methods such as starspot tracking or gravity darkening, Siegel et al. 2023; Dong & Foreman-Mackey 2023) and constitute the majority of obliquity measurements (Triaud, 2018; Albrecht et al., 2022), prioritizing those selected in Albrecht et al. (2022) (we otherwise adopt the “preferred” measurements from TEPCat). Finally, we apply a mass cut of , and select only single-star systems by cross-matching our sample with the multi-star catalog of Rice et al. (2024), which is based on the most recent data from Gaia DR3. We distinguish between sub-Saturns that are isolated and those in compact multi-planet systems, which we define as having a nearby companion with a period ratio of . All sub-Saturns considered in this work are presented in Table 2 (note that we supplement any missing parameters using the NASA Exoplanet Archive555https://exoplanetarchive.ipac.caltech.edu and values adopted by their respective reference papers). Our RM measurements allow for more reliable statistical analyses to be performed on the relationship between compact configurations and spin-orbit alignment for sub-Saturns (Section 5.1) and thus more robust constraints on the origins of spin-orbit misalignment as well as the formation histories of the sub-Saturn population (Sections 5.2 and 5.3).
| System | Planet | (K) | [] | (d) | () | () | Reference |
| Compact Multi-planet Systems: | |||||||
| AU Mic | b | 3665 | 0.5 | 8.46 | 20.12 | Hirano et al. (2020a) | |
| Kepler-9 | b | 5774 | 1.02 | 19.24 | 43.41 | 13 | Wang et al. (2018) |
| WASP-148 | b | 5555 | 0.95 | 8.80 | 91.24 | Wang et al. (2022) | |
| TOI-5126 | b | 6253 | 1.26 | 5.46 | 21.3 | 1 | this work |
| TOI-5398 | b | 6028 | 1.105 | 10.59 | 54.4 | this work | |
| Isolated Systems: | |||||||
| GJ 436 | b | 3586 | 0.47 | 2.64 | 22.25 | 72 | Bourrier et al. (2018) |
| HAT-P-11 | b | 4780 | 0.81 | 4.89 | 26.7 | 103 | Winn et al. (2010b) |
| HAT-P-12 | b | 4650 | 0.73 | 3.21 | 67.08 | Mancini et al. (2018) | |
| HD 89345 | b | 5576 | 1.16 | 11.81 | 34.97 | 74.2 | Bourrier et al. (2023) |
| HD 332231 | b | 6089 | 1.13 | 18.71 | 77.57 | Knudstrup & Albrecht (2022) | |
| HIP 67522 | b | 5675 | 1.22 | 6.96 | Heitzmann et al. (2021) | ||
| K2-25 | b | 3207 | 0.26 | 3.48 | 24.51 | 3 | Stefansson et al. (2020) |
| K2-105 | b | 5636 | 1.05 | 8.27 | 30.01 | Bourrier et al. (2023) | |
| KELT-11 | b | 5375 | 1.44 | 4.74 | 54.36 | Mounzer et al. (2022) | |
| TOI-1842 | b | 6230 | 1.46 | 9.57 | 68.03 | Hixenbaugh et al. (2023) | |
| TOI-1859 | b | 6341 | 1.29 | 63.48 | 38.9 | Dong et al. (2023) | |
| WASP-39 | b | 5400 | 0.93 | 4.06 | 89.01 | 0 | Mancini et al. (2018) |
| WASP-69 | b | 4700 | 0.98 | 3.87 | 92.19 | 0.4 | Casasayas-Barris et al. (2017) |
| WASP-107 | b | 4425 | 0.68 | 5.72 | 30.52 | 118.1 | Rubenzahl et al. (2021) |
| WASP-131 | b | 6030 | 1.06 | 5.32 | 85.83 | 162.4 | Doyle et al. (2023) |
| WASP-156 | b | 4910 | 0.84 | 3.84 | 40.69 | Bourrier et al. (2023) | |
| WASP-166 | b | 6050 | 1.19 | 5.44 | 32.11 | 3 | Hellier et al. (2019) |
5.1 Sub-Saturns in Compact Multi-planet Systems are Significantly Aligned
It is clear from Figure 2 that spin-orbit alignment is much more common for sub-Saturns in compact multi-planet systems than for isolated sub-Saturns, while no clear relation exists for the host star’s effective temperature (as seen in previous works, e.g., Albrecht et al. 2022). Our two stellar obliquity measurements increase the sample of sub-Saturns in compact configurations to five, enabling the first statistical tests of this trend. We investigate the statistical significance of preferential alignment in compact systems by comparing the number of misaligned systems found in the compact multi-planet sub-Saturn sample (zero) with that found in a series of random draws from the isolated sub-Saturn sample.
Namely, of the 22 sub-Saturns with measurements considered in this work, five are in compact multi-planet systems, while the other 17 are isolated. Of the five in compact configurations, none are misaligned, while 9/17 (53%) isolated sub-Saturns are misaligned. We define misalignment as at the level (i.e., ) and at the level (i.e., ), where we adopt the uncertainties () reported for each measurement (see Table 2). In order to perform a fair comparison between the compact multi-planet sample and the isolated sample, we randomly select and classify the alignment of five isolated sub-Saturns. We perform 100,000 iterations of these random draws, counting the number of misaligned systems for each, and fit a Gaussian function over the resulting distribution. The upper-right sub-panel of Figure 2 reveals the result, illustrating that sub-Saturns in compact multi-planet systems are more often aligned than isolated sub-Saturns to a significance level of .
We verified that this significance is robust to the precise misalignment criteria adopted by re-running the above procedure with the following alternative cuts: and at the level ( significance), (ii) and at the level ( significance), (iii) ( significance), and (iv) ( significance). Assuming our original alignment criteria, we additionally vary the Neptune/sub-Saturn mass boundary and re-run our procedure: (i) ( significance), and (ii) ( significance). We note that while the significance appears to degrade with an increasing mass bound, this trend is likely not astrophysical, as the significance is directly proportional to the number of sub-Saturns in compact systems considered, which drops from six assuming a bound to only three assuming a () bound. Finally, we note that the young (), aligned, compact sub-Saturn system AU Mic still hosts a dusty debris disk (Plavchan et al., 2020; Hirano et al., 2020a), meaning that its spin-orbit angle and compactness has likely not yet had the opportunity to be affected by various pathways to trigger post-disk eccentric migration (Wu et al., 2023). Therefore, while this system provides strong support for primordial alignment, it may bias our statistical test; we re-run our original procedure excluding AU Mic (such that there are only four compact sub-Saturn systems) and find that the significance is maintained, but reduced to .
5.2 Evidence for Primordial Alignment and Misalignment Via Eccentric Migration
In this work, we have demonstrated that sub-Saturns in compact planetary systems are significantly more aligned than those without nearby planetary companions (at the level). This trend of spin-orbit alignment in compact multi-planet systems around single stars is not limited to sub-Saturns, however. Short-period () Jupiters provided the first robust evidence for this phenomenon. Specifically, all measured spin-orbit angles for short-period Jupiters with nearby companions, even hot Jupiters, are consistent with alignment, e.g., Kepler-30 (Sanchis-Ojeda et al., 2012), Kepler-89 (Albrecht et al., 2013), WASP-47 (Sanchis-Ojeda et al., 2015), WASP-148 (Wang et al., 2022), TOI-1478 Rice et al. (2022), TOI-2202 (Rice et al., 2023b), TOI-1670 (Lubin et al., 2023). Conversely, isolated short-period Jupiters display a wide range of obliquities and are frequently misaligned (Albrecht et al. 2012, 2022). While there is a paucity of spin-orbit measurements for sub-Neptunes and super-Earths, most of these lower-mass planets reside in compact multi-planet systems, and those with secure obliquity measurements are aligned, e.g., Kepler-50, Kepler-65 (Chaplin et al., 2013), Kepler-25 (Albrecht et al., 2013), HD 106315 (Zhou et al., 2018; Bourrier et al., 2023), TRAPPIST-1 (Hirano et al., 2020b), TOI-1726 (Dai et al., 2020), HD 63433 (Mann et al., 2020), TOI-1136 (Dai et al., 2023), and TOI-2076 (Frazier et al., 2023).
As the spin-orbit angles of compact multi-planet systems should remain largely unaltered following the dispersal of the disk, the low stellar obliquities observed for sub-Saturns and other types of exoplanets in compact systems suggest that planetary systems are initially formed spin-orbit aligned. Furthermore, the large misalignments observed for isolated hot Jupiters and sub-Saturns are generally consistent with violent eccentric migration pathways (Dawson & Johnson, 2018; Wu et al., 2023). This scenario is additionally supported by observational evidence showing that i) both young stellar systems ( old) and planets still embedded within their debris disks are aligned (Kraus et al., 2020; Plavchan et al., 2020; Hirano et al., 2020b; Martioli et al., 2020; Palle et al., 2020; Carmichael et al., 2021; Albrecht et al., 2022; Johnson et al., 2022), and ii) that misaligned hot Jupiter systems tend to be older (Hamer & Schlaufman, 2022), reinforcing the notion that misalignments are acquired well into in the post-disk phase rather than near the onset of system formation.
5.3 On the Temperature-Obliquity Relation
There is an interesting discrepancy between the obliquity distributions of isolated hot Jupiters and isolated hot sub-Saturns: hot Jupiters are commonly misaligned around hot stars but exclusively aligned around cool stars (i.e., the relation, see Winn et al. 2010a; Albrecht et al. 2012, 2022), while hot sub-Saturns are commonly misaligned around cool stars (and no data exist for those orbiting hot stars, see Figure 2). Realignment via tidal dissipation offers a plausible explanation for this phenomenon; cool stars are susceptible to realignment by massive hot Jupiters, while lower-mass hot sub-Saturns are less likely to realign them on timescales shorter than the system’s age (e.g., see Eqn. 2 of Albrecht et al. 2012). Hence, both hot Jupiters and hot sub-Saturns may indeed become misaligned through high-eccentricity migration, but only cool stars with hot Jupiters can be realigned. This tidal realignment mechanism, however, requires finely-tuned parameters in order to completely realign all hot Jupiters around cool stars without leaving any on polar or retrograde orbits (Rogers et al., 2013; Li & Winn, 2016), which are not seen in observations.
Alternatively or in addition to this tidal realignment scenario, it is possible that sub-Saturns are subject to misalignment mechanisms that may not operate for Jupiters. For instance, Petrovich et al. (2020) illustrates that secular resonance between an outer gas giant companion and an inner sub-Saturn, both embedded in a decaying disk, can drive the sub-Saturn to a polar orbit. The level of obliquity excitation depends on the ratio of the angular momentum between the inner and outer orbits, with nearly polar orbits preferentially excited for inner sub-Saturn-mass planets over Jupiter-mass planets. This resonance sweeping offers a natural explanation for the excess of sub-Saturns observed to be on polar orbits, regardless of stellar effective temperature (Bourrier et al., 2018; Hixenbaugh et al., 2023), as well as a potential mechanism to stunt the runaway growth of sub-Saturns since they are assumed to be embedded within an actively dispersing disk. However, this mechanism predicts similar misalignment rates for compact multi-planet and isolated systems, holding all other factors equal. This would imply that isolated sub-Saturns, which are more often misaligned, host outer companions at a higher rate than sub-Saturns in compact systems, but no statistical studies on these relative rates have yet been completed.
To explain the relation and spin-orbit misalignment in single-star exoplanetary systems more broadly, Hixenbaugh et al. (2023) outline a unified model wherein planets gain misalignments if other planets of comparable (or higher) mass form within their disk. Specifically, hotter, more massive stars tend to host more massive protoplanetary disks (Williams & Cieza, 2011; Andrews et al., 2013; Andrews, 2020), which are more likely to form multiple Jupiter-mass planets (Johnson et al., 2010; Ghezzi et al., 2018; Yang et al., 2020). These Jupiters may interact with each other to produce spin-orbit misalignments after the disk dissipates. By contrast, the limited disk mass around cooler stars is unlikely to produce more than one Jupiter-mass planet, if any (Andrews et al., 2013; Ansdell et al., 2016; Pascucci et al., 2016; Dawson & Johnson, 2018; Yang et al., 2020). As a result, these giants, which have no comparable-mass planetary perturbers, should largely retain their primordial alignments throughout the post-disk phase. Nevertheless, despite their less massive disks, cool stars may still form several sub-Saturn planets that undergo an array of dynamical interactions in the post-disk phase that are capable of exciting their stellar obliquities, analogously to multiple Jupiters around hot stars. As present-day compact systems may have avoided violent dynamical interactions, they preserve their primordial spin-orbit angle, allowing them to serve as a window into the early, undisturbed state of planetary systems.
Acknowledgments
We acknowledge the independent work on the Rossiter-McLaughlin effect measurement of TOI-5398 b by Mantovan et al. Two manuscripts were prepared concurrently but independently. We appreciate the collegiality of the team regarding the manuscript submission process.
We express our gratitude for the insightful discussions with George Zhou and Chelsea Huang regarding the stellar activity of young planetary systems. We would also like to thank Armaan Goyal for insightful discussions regarding statistical analyses.
Data presented were obtained by the NEID spectrograph built by Penn State University and operated at the WIYN Observatory by NOIRLab, under the NN-EXPLORE partnership of the National Aeronautics and Space Administration and the National Science Foundation. These results are based on observations obtained with NEID on the WIYN 3.5m Telescope at Kitt Peak National Observatory (PI: Songhu Wang, IU TAC). Additional NEID data were obtained through time allocated by NSF’s NOIRLab (NOIRLab Prop. IDs 2022A-543544, 2023A-904842, 2023B-738235; PI: S. Yee), under the NN-EXPLORE partnership of the National Aeronautics and Space Administration and the National Science Foundation. WIYN is a joint facility of the University of Wisconsin–Madison, Indiana University, NSF’s NOIRLab, the Pennsylvania State University, Purdue University, University of California, Irvine, and the University of Missouri. The authors are honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham.
M.R. acknowledges support from Heising-Simons Foundation grant 2023-4478. S.W. acknowledges support from Heising-Simons Foundation grant 2023- 4050. We acknowledge support from the NASA Exoplanets Research Program NNH23ZDA001N-XRP (grant 80NSSC24K0153). This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.
All the TESS data used in this paper can be found in MAST: doi:10.17909/t9-nmc8-f686 (catalog 10.17909/t9-nmc8-f686). This research has made use of the NASA Exoplanet Archive (NASA Exoplanet Archive, 2024), which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.